# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 3.2.1 Staying in a nice set of configurations
We introduce some notations that will be useful in what follows. Recall for $\lambda>1$, that $\alpha(\lambda)$ is defined to be the largest positive root of the equation (7) (i.e., $\alpha(\lambda)n$ is approximately the expected size of the giant in a $G(n,d/n)$), and recall $\sigma^{2}(\lambda)$ as defined in (9) which approximates the variance of the giant component in $G(n,\lambda/n)$. Also, for $\theta\in[0,1]$, let $k_{\mathsf{a}}(\theta)$ be the expected fraction of activated vertices if a giant of fractional size $\theta$ is activated, and $k_{\mathsf{ia}}(\theta)$ be the same quantity if it is not activated, i.e.,
@@
Let $\beta=\beta_{\mathsf{c}}$. For all $T\geq 0$ fixed independent of $n$, if  $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$, then for $K=K(T)$, $X_{T}\in\mathcal{G}_{K}$ with probability $1-o(1)$.
@@
Let ${X}_{0}\sim G(n,\frac{p_{0}}{n})$. For some $K_{0}>0$, Lemmas 2.11 implies that, ${X}_{0}\in\mathcal{G}_{K_{0}}$ with $1-o(1)$ probability. Let $K_{t}=tK_{0}$. Now suppose ${X}_{t}\in\mathcal{G}_{K_{t}}$ for some $t\geq 0$. We will inductively show that ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$ with probability $1-o(1)$. Then a union bound over $t\in[0,T]$ implies ${X}_{T}\in\mathcal{G}_{K_{T}}$.
@@
where $C_{1}>0$ is some constant. In the percolation step, $G_{t+1}\sim G(A_{t+1},\beta_{\mathsf{c}}/n)$, and when the concentration in (26) holds, the random graph $G_{t+1}$ is supercritical. By Lemma 2.10, $|\mathcal{L}_{2}(G_{t+1})|=O(\log n)$ with probability $1-O(n^{-1})$ and $||\mathcal{L}_{1}(G_{t+1})|-\theta_{*}n|\leq n^{1/2}\log n$ with probability $1-o(n^{-1})$. Since $|\mathcal{L}_{2}(X_{t})|\leq K_{t}\log n$ by property 2 and the largest component of ${X}_{t}$ has been activated, if $|\mathcal{L}_{1}(G_{t+1})|=\Omega(n)$, then $\mathcal{L}_{1}({X}_{t+1})=\mathcal{L}_{1}(G_{t+1})$. Moreover, with probability $1-o(1)$, we have $G_{t+1}\in\mathcal{G}_{K_{0}}$ by Lemma 2.11. Consequently, we also establish property 2-5 by noting that
@@
where $C_{2}>0$ is a constant. In the percolation step, $G_{t+1}\sim G(A_{t+1},\beta_{\mathsf{c}}/n)$ and if the estimate in (27) holds then $G_{t+1}$ is subcritical. In this case, by Lemma 2.8, $|\mathcal{L}_{1}(G_{t+1})|\leq K_{0}\log n$ with probability $1-o(1)$, so $L_{t+1}=L_{t}$, satisfying the first property of $\mathcal{G}_{(t+1)K_{0}}$. In addition, with probability $1-o(1)$, $G_{t+1}$ satisfies that $R_{2}(G_{t+1})\leq K_{0}n,R_{3}(G_{t+1})\leq K_{0}n,I_{1}(G_{t+1})\geq n/K_{0}$. Similar to the case $\Lambda_{t+1}$, these properties combined with the hypothesis ${X}_{t}\in\mathcal{G}_{K_{t}}$ imply that ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$. Therefore, by a union bound over all the above, we have ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$ with probability $1-o(1)$. ∎
@@
Let $\beta=\beta_{\mathsf{c}}$. For all $T\geq 0$ fixed independent of $n$, if  $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$, then $(X_{s})_{s\leq T}\in\mathcal{G}_{T}$, with probability $1-o(1)$.
@@
Fix $T\geq 0$ We will show that for each $s\leq T$, the event in Definition 3.13 holds with probability $1-o(1)$; the lemma follows from a union bound over all $s\leq T$. First, Lemma 3.12 shows that $X_{0},...,X_{T}\in\mathcal{G}_{K}$ with probability $1-o(1)$. Working on this event, when $X_{0},\dots,X_{s}\in\mathcal{G}_{K}$ we can express
@@
where we are using $G_{i}^{t}$ to denote the (random) subgraph induced on the vertex set of $G_{i}$ (the resampled portion of $X_{i}$), consisting of vertices that are not re-activated up through time $t$ (i.e., do not belong to $\bigcup_{j=i+1}^{t}G_{j}$). That is, the vertices of $G_{i}$ that survive $t-i$ independent activation steps. Eq. (29) is a consequence of the fact that $X_{t}$ is partitioned by $\bigcup_{i=0}^{t}G_{i}^{t}$ ($G_{t}^{t}$ is the activated subgraph at time $t$).
@@
By (29) and a union bound of (30)–(31) over $i=1,\dots,s$, together with Observation 3.25, we get the claimed bound for $R_{2}^{-}(X_{s})$ with probability $1-o(1)$.
@@
By (26), we have $|A_{i}-k_{\mathsf{a}}(L_{i-1}/n)n|\leq\sqrt{n}\log n$ with probability $1-o(1/n)$; by the first property of $\mathcal{G}_{K}$, we also have $|k_{\mathsf{a}}(L_{i-1}/n)n-k_{\mathsf{a}}(\theta_{*})n|\leq\sqrt{n}\log n$. Let us work on the event that $A_{i}\in W_{\mathsf{a}}$. We will rely on the following lemma which establishes concentration of the number of vertices not yet activated after $s-i$ steps; its proof follows from a simple calculation and Chebyshev’s inequality, and is deferred.
@@
$\mathbb{P}\big{(}\big{|}R_{2}^{-}(G_{i}^{s})-\mathbb{E}[R_{2}(G_{*}^{s-i})]\big{|}=O(\sqrt{n}\log n)\,,\,A_{i}\in W_{\mathsf{a}}\big{)}=1-O\big{(}\frac{1}{\log^{2}n}\big{)}\,.$
Next suppose that that $\Lambda_{i}$ does not happen. Set $W_{\mathsf{ia}}:=[nk_{\mathsf{ia}}(\theta_{*})-\sqrt{n}\log n,nk_{\mathsf{ia}}(\theta_{*})+\sqrt{n}\log n]$. By (27) we have $A_{i}\in W_{\mathsf{ia}}$ with probability $1-o(1/n)$. In this case, $G(A_{i},\frac{\beta_{\mathsf{c}}}{n})$ is subcritical (uniformly in $n$). Combining Observation 3.25 and (12) in Lemma 2.7, we get that
@@
$\mathbb{P}\big{(}\big{|}R_{2}(G_{i}^{s})-\mathbb{E}[R_{2}(G_{*}^{s-i})]\big{|}=O(\sqrt{n}\log n)\,,\,A_{i}\in W_{\mathsf{ia}}\big{)}=O\big{(}\frac{1}{\log^{2}n}\big{)}.$ (32)
Hence, we get (30) by a union bound. The argument for (31) is only easier than the above, following immediately from Lemma 3.15 and the super-criticality of the initialization parameter. ∎
@@ 3.2.2 Upper bound on escape time from the unstable fixed point
If $L_{0}\notin I_{\gamma}$, then $\tilde{\tau}=0$ and the lemma holds trivially. We assume therefore that $L_{0}\in I_{\gamma}$. Also we assume $X_{0}\in\mathcal{G}_{K}$, for a large constant $K$, which happens with probability $1-o(1)$ by Lemma 3.12. Let $\tau^{\prime}:=\min\{t:X_{t}\notin\mathcal{G}_{K}\}$ and let $\tau=\tilde{\tau}\wedge\tau^{\prime}$. Observe that $\tau$ is stochastically dominated from above by a geometric random variable $\text{Geo}(y)$, where
@@
Indeed, assuming (33), $\mathbb{E}[\tau\mid X_{0}\in I_{\gamma}]\leq\mathbb{E}[\text{Geo}(y)]\leq e^{\kappa\gamma^{2}}$, and by Markov’s inequality, we would have $\mathbb{P}(\tau>2Te^{\kappa\gamma^{2}})\leq\frac{1}{2T}$ for every $T$. Then we have
@@
Suppose $L(X_{t})=x_{0}\in I_{\gamma}$. By Chebyshev’s inequality we have
@@ 3.2.3 Approximating by a 1-dimensional Markov process
Suppose $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$ and let $(Z_{t})_{t}$ be the process of Definition 3.18 initialized from $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$. There exists a coupling such that with probability $1-T\cdot o(1)$, for all $t\leq T$,
@@
We will couple $(L_{t},Y_{t})$ inductively. For the base case, consider $t=0$; by Lemma 2.3, $(X_{0},Y_{0})$ can be coupled such that with probability $1-o(1)$, $|L_{0}-Y_{0}|\leq 1$.
Now suppose there exists a coupling of the first $t$ steps of $\{({X}_{t},Y_{t})\}$ such that $|L_{t}-Y_{t}|\leq 1$ with probability $1-o(1)$. By Lemma 3.14, we have $({X}_{s})_{s\leq t}\in\mathcal{G}_{t}$ with probability $1-o(1)$, so in what follows we work on that event. In case $({X}_{s})_{s\leq t}\notin\mathcal{G}_{t}$, we stop the coupling. For the induction step, we construct a coupling $\operatorname{{\mathbf{P}}}_{t+1}$ of $({X}_{t+1},Y_{t+1})$.
On $\Lambda^{c}_{t+1}$, we let $Y_{t+1}=Y_{t}$. At the same time, on $\Lambda^{c}_{t+1}$ and $\mathcal{G}_{t}$, with high probability the percolation step is sub-critical and therefore $L_{t+1}=L_{t}$ with probability $1-o(1)$ (a similar argument was made in the proof of Lemma 3.12). Hence, we have $|L_{t+1}-Y_{t+1}|=|L_{t}-Y_{t}|\leq 1$ with probability $1-o(1)$ in this case.
@@
$\|B_{t+1}-{N}_{t+1}\|_{\textsc{tv}}\leq\frac{3|\text{Var}(A_{t+1}^{-}\mid{X}_{t})-\sigma_{t}^{2}|}{2\text{Var}(A^{-}_{t+1}\mid{X}_{t})}+\frac{\left|\mathbb{E}[A^{-}_{t+1}\mid{X}_{t}]-\mathbb{E}[B_{t+1}\mid Y_{t}]\right|}{2\sqrt{\text{Var}(A^{-}_{t+1}\mid{X}_{t})}}=O\left(\frac{\sqrt{n}\log^{2}n}{n}\right)=O\left(\frac{\log^{2}n}{\sqrt{n}}\right).$
@@
On the percolation step, Lemma 3.29 implies that there exists a coupling of $(|\mathcal{L}_{1}(G_{t+1})|,Y_{t+1})$ such that $|Y_{t+1}-|\mathcal{L}_{1}(G_{t+1})||\leq 1$ with probability $1-o(1)$. Since $|\mathcal{L}_{2}({X}_{t})|=O(\log n)$ and the giant component has been activated, with high probability the percolation step is super-critical and, $L_{t+1}=|\mathcal{L}_{1}(G_{t+1})|$.
Therefore, it follows a union bound over all the probabilistic estimates used so far that the coupling of $({X}_{t+1},Y_{t+1})$ satisfies all the desired properties with probability $1-o(1)$. ∎
@@
$=\theta_{*}n+(1+f^{\prime}(\theta_{*}))(\Theta_{t}-\theta_{*})n+O(\gamma^{2}).$
@@
Finally we handle the normal random variable in (42). Let $\tilde{d}_{*}:=\beta_{\mathsf{c}}\tilde{k}(\theta_{*})$ and $\tilde{d}_{t}:=\beta_{\mathsf{c}}\cdot\tilde{k}(\Theta_{t})$. Since $h_{2}^{\prime}$ is a bounded function and $|\Theta_{t}-\theta_{*}|\leq\frac{\gamma}{\sqrt{n}}$, $|h_{2}(\tilde{d}_{t})-h_{2}(\tilde{d}_{*})|=O(\frac{\gamma}{\sqrt{n}})$. Again, since $h_{2}^{\prime}$ is a bounded function and $|\tilde{d}_{*}-d_{*}|=O\big{(}\sqrt{\frac{\log n}{n}}\big{)}$, $|h_{2}(\tilde{d}_{*})-h_{2}(d_{*})|=O\big{(}\sqrt{\frac{\log n}{n}}\big{)}.$ By triangle inequality,

### v2

@@ 3.2.1 Staying in a nice set of configurations
We introduce some notations that will be useful in what follows. Recall for $\lambda>1$, that $\alpha(\lambda)$ is defined to be the largest positive root of the equation (7) (i.e., $\alpha(\lambda)n$ is approximately the expected size of the giant in a $G(n,\lambda/n)$), and recall $\sigma^{2}(\lambda)$ as defined in (9) which approximates the variance of the giant component in $G(n,\lambda/n)$. Also, for $\theta\in[0,1]$, let $k_{\mathsf{a}}(\theta)$ be the expected fraction of activated vertices if a giant of fractional size $\theta$ is activated, and $k_{\mathsf{ia}}(\theta)$ be the same quantity if it is not activated, i.e.,
@@
Let $\beta=\beta_{\mathsf{c}}$. For all $T\geq 0$ fixed independent of $n$, if  $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$, then for $K=K(T)$, $X_{T}\in\mathcal{G}_{K}$ with probability 1o(1)1-o(1).
@@
Let ${X}_{0}\sim G(n,\frac{p_{0}}{n})$. For some $K_{0}>0$, Lemma 2.11 implies that, ${X}_{0}\in\mathcal{G}_{K_{0}}$ with 1o(1)1-o(1) probability. Let $K_{t}=tK_{0}$. Now suppose ${X}_{t}\in\mathcal{G}_{K_{t}}$ for some $t\geq 0$. We will inductively show that ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$ with probability 1o(1)1-o(1). Then a union bound over $t\in[0,T]$ implies ${X}_{T}\in\mathcal{G}_{K_{T}}$.
@@
where $C_{1}>0$ is some constant. In the percolation step, $G_{t+1}\sim G(A_{t+1},\beta_{\mathsf{c}}/n)$, and when the concentration in (26) holds, the random graph $G_{t+1}$ is supercritical. By Lemma 2.10, $|\mathcal{L}_{2}(G_{t+1})|=O(\log n)$ with probability $1-O(n^{-1})$ and $||\mathcal{L}_{1}(G_{t+1})|-\theta_{*}n|\leq n^{1/2}\log n$ with probability $1-o(n^{-1})$. Since $|\mathcal{L}_{2}(X_{t})|\leq K_{t}\log n$ by property 2 and the largest component of ${X}_{t}$ has been activated, if $|\mathcal{L}_{1}(G_{t+1})|=\Omega(n)$, then $\mathcal{L}_{1}({X}_{t+1})=\mathcal{L}_{1}(G_{t+1})$. Moreover, with probability 1o(1)1-o(1), we have $G_{t+1}\in\mathcal{G}_{K_{0}}$ by Lemma 2.11. Consequently, we also establish property 2-5 by noting that
@@
where $C_{2}>0$ is a constant. In the percolation step, $G_{t+1}\sim G(A_{t+1},\beta_{\mathsf{c}}/n)$ and if the estimate in (27) holds then $G_{t+1}$ is subcritical. In this case, by Lemma 2.8, $|\mathcal{L}_{1}(G_{t+1})|\leq K_{0}\log n$ with probability 1o(1)1-o(1), so $L_{t+1}=L_{t}$, satisfying the first property of $\mathcal{G}_{(t+1)K_{0}}$. In addition, with probability 1o(1)1-o(1), $G_{t+1}$ satisfies that $R_{2}(G_{t+1})\leq K_{0}n,R_{3}(G_{t+1})\leq K_{0}n,I_{1}(G_{t+1})\geq n/K_{0}$. Similar to the case $\Lambda_{t+1}$, these properties combined with the hypothesis ${X}_{t}\in\mathcal{G}_{K_{t}}$ imply that ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$. Therefore, by a union bound over all the above, we have ${X}_{t+1}\in\mathcal{G}_{K_{t+1}}$ with probability 1o(1)1-o(1). ∎
@@
Let $\beta=\beta_{\mathsf{c}}$. For all $T\geq 0$ fixed independent of $n$, if  $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$, then $(X_{s})_{s\leq T}\in\mathcal{G}_{T}$, with probability 1o(1)1-o(1).
@@
Fix $T\geq 0$. We will show that for each $s\leq T$, the event in Definition 3.13 holds with probability 1o(1)1-o(1); the lemma follows from a union bound over all $s\leq T$. First, Lemma 3.12 shows that $X_{0},...,X_{T}\in\mathcal{G}_{K}$ with probability 1o(1)1-o(1). Working on this event, when $X_{0},\dots,X_{s}\in\mathcal{G}_{K}$ we can express
@@
where we are using $G_{i}^{t}$ to denote the (random) subgraph induced on the vertex set of $G_{i}$ (the resampled portion of $X_{i}$), consisting of vertices that are not re-activated up through time $t$ (i.e., do not belong to $\bigcup_{j=i+1}^{t}G_{j}$). That is, the vertices of $G_{i}$ that survive tit-i independent activation steps. Eq. (29) is a consequence of the fact that $X_{t}$ is partitioned by $\bigcup_{i=0}^{t}G_{i}^{t}$ ($G_{t}^{t}$ is the activated subgraph at time $t$).
@@
By (29) and a union bound of (30)–(31) over $i=1,\dots,s$, together with Observation 3.25, we get the claimed bound for $R_{2}^{-}(X_{s})$ with probability 1o(1)1-o(1).
@@
By (26), we have $|A_{i}-k_{\mathsf{a}}(L_{i-1}/n)n|\leq\sqrt{n}\log n$ with probability 1o(1/n)1-o(1/n); by the first property of $\mathcal{G}_{K}$, we also have $|k_{\mathsf{a}}(L_{i-1}/n)n-k_{\mathsf{a}}(\theta_{*})n|\leq\sqrt{n}\log n$. Let us work on the event that $A_{i}\in W_{\mathsf{a}}$. We will rely on the following lemma which establishes concentration of the number of vertices not yet activated after sis-i steps; its proof follows from a simple calculation and Chebyshev’s inequality, and is deferred.
@@
$\mathbb{P}\big{(}\big{|}R_{2}^{-}(G_{i}^{s})-\mathbb{E}[R_{2}(G_{*}^{s-i})]\big{|}=O(\sqrt{n}\log n)\mid A_{i}\in W_{\mathsf{a}}\big{)}=1-O\big{(}\frac{1}{\log^{2}n}\big{)}\,.$
Next suppose that that $\Lambda_{i}$ does not happen. Set $W_{\mathsf{ia}}:=[nk_{\mathsf{ia}}(\theta_{*})-\sqrt{n}\log n,nk_{\mathsf{ia}}(\theta_{*})+\sqrt{n}\log n]$. By (27) we have $A_{i}\in W_{\mathsf{ia}}$ with probability 1o(1/n)1-o(1/n). In this case, $G(A_{i},\frac{\beta_{\mathsf{c}}}{n})$ is subcritical (uniformly in $n$). Combining Observation 3.25 and (12) in Lemma 2.7, we get that
@@
$\mathbb{P}\big{(}\big{|}R_{2}(G_{i}^{s})-\mathbb{E}[R_{2}(G_{*}^{s-i})]\big{|}=O(\sqrt{n}\log n)\mid A_{i}\in W_{\mathsf{ia}}\big{)}=1-O\big{(}\frac{1}{\log^{2}n}\big{)}.$ (32)
Hence, we get (30) by conditioning on $\Lambda_{i}$ and its complement and a union bound. The argument for (31) is only easier than the above, following immediately from Lemma 3.15 and the super-criticality of the initialization parameter. ∎
@@ 3.2.2 Upper bound on escape time from the unstable fixed point
If $L_{0}\notin I_{\gamma}$, then $\tilde{\tau}=0$ and the lemma holds trivially. We assume therefore that $L_{0}\in I_{\gamma}$. Also we assume $X_{0}\in\mathcal{G}_{K}$, for a large constant $K$, which happens with probability 1o(1)1-o(1) by Lemma 3.12. Let $\tau^{\prime}:=\min\{t:X_{t}\notin\mathcal{G}_{K}\}$ and let $\tau=\tilde{\tau}\wedge\tau^{\prime}$. Observe that $\tau$ is stochastically dominated from above by a geometric random variable $\text{Geo}(y_{\gamma})$, where
@@
Indeed, assuming (33), $\mathbb{E}[\tau\mid X_{0}\in I_{\gamma}]\leq\mathbb{E}[\text{Geo}(y_{\gamma})]\leq e^{\kappa\gamma^{2}}$, and by Markov’s inequality, we would have $\mathbb{P}(\tau>2Te^{\kappa\gamma^{2}})\leq\frac{1}{2T}$ for every $T$. Then we have
@@
Suppose $L_{t}=x_{0}\in I_{\gamma}$. By Chebyshev’s inequality we have
@@ 3.2.3 Approximating by a 1-dimensional Markov process
Suppose $X_{0}\sim G(n,\frac{p_{0}}{n})$ with $p_{0}=\lambda_{*}+O(n^{-1/2})$ and let $(Z_{t})_{t}$ be the process of Definition 3.18 initialized from $Z_{0}\sim\mathcal{N}((p_{0}-\lambda_{*})\alpha^{\prime}(\lambda_{*})n,\sigma^{2}(\lambda_{*})n)$. For all $T\geq 0$ fixed independent of $n$, there exists a coupling such that with probability 1o(1)1-o(1), for all $t\leq T$,
@@
We will couple $(L_{t},Y_{t})$ inductively. For the base case, consider $t=0$; by Lemma 2.3, $(X_{0},Y_{0})$ can be coupled such that with probability 1o(1)1-o(1), $|L_{0}-Y_{0}|\leq 1$.
Now suppose there exists a coupling of the first $t$ steps of $\{({X}_{t},Y_{t})\}$ such that $|L_{t}-Y_{t}|\leq 1$ with probability 1o(1)1-o(1). By Lemma 3.14, we have $({X}_{s})_{s\leq t}\in\mathcal{G}_{t}$ with probability 1o(1)1-o(1), so in what follows we work on that event. In case $({X}_{s})_{s\leq t}\notin\mathcal{G}_{t}$, we stop the coupling. For the induction step, we construct a coupling $\operatorname{{\mathbf{P}}}_{t+1}$ of $({X}_{t+1},Y_{t+1})$.
On $\Lambda^{c}_{t+1}$, we let $Y_{t+1}=Y_{t}$. At the same time, on $\Lambda^{c}_{t+1}$ and $\mathcal{G}_{t}$, with high probability the percolation step is sub-critical and therefore $L_{t+1}=L_{t}$ with probability 1o(1)1-o(1) (a similar argument was made in the proof of Lemma 3.12). Hence, we have $|L_{t+1}-Y_{t+1}|=|L_{t}-Y_{t}|\leq 1$ with probability 1o(1)1-o(1) in this case.
@@
$\|B_{t+1}-{N}_{t+1}\|_{\textsc{tv}}$ $\leq\frac{3|\text{Var}(A_{t+1}^{-}\mid{X}_{t})-\sigma_{t}^{2}|}{2\text{Var}(A^{-}_{t+1}\mid{X}_{t})}+\frac{\left|\mathbb{E}[A^{-}_{t+1}\mid{X}_{t}]-\mathbb{E}[B_{t+1}\mid Y_{t}]\right|}{2\sqrt{\text{Var}(A^{-}_{t+1}\mid{X}_{t})}}$
$=O\left(\frac{\sqrt{n}\log^{2}n}{n}\right)=O\left(\frac{\log^{2}n}{\sqrt{n}}\right).$
@@
On the percolation step, Lemma 3.29 implies that there exists a coupling of $(|\mathcal{L}_{1}(G_{t+1})|,Y_{t+1})$ such that $|Y_{t+1}-|\mathcal{L}_{1}(G_{t+1})||\leq 1$ with probability 1o(1)1-o(1). Since $|\mathcal{L}_{2}({X}_{t})|=O(\log n)$ and the giant component has been activated, with high probability the percolation step is super-critical and, $L_{t+1}=|\mathcal{L}_{1}(G_{t+1})|$.
Therefore, it follows a union bound over all the probabilistic estimates used so far that the coupling of $({X}_{t+1},Y_{t+1})$ satisfies all the desired properties with probability 1o(1)1-o(1). ∎
@@
$=\theta_{*}n+(1+f^{\prime}(\theta_{*}))(\Theta_{t}-\theta_{*})n+O(1).$
@@
Finally we handle the normal random variable in (42). Let $\tilde{d}_{*}:=\beta_{\mathsf{c}}\tilde{k}(\theta_{*})$ and $\tilde{d}_{t}:=\beta_{\mathsf{c}}\cdot\tilde{k}(\Theta_{t})$. Since $h_{2}^{\prime}$ is a bounded function and $|\Theta_{t}-\theta_{*}|\leq\frac{\gamma}{\sqrt{n}}$, $|h_{2}(\tilde{d}_{t})-h_{2}(\tilde{d}_{*})|=O(\frac{1}{\sqrt{n}})$. Again, since $h_{2}^{\prime}$ is a bounded function and $|\tilde{d}_{*}-d_{*}|=O\big{(}\sqrt{\frac{\log n}{n}}\big{)}$, $|h_{2}(\tilde{d}_{*})-h_{2}(d_{*})|=O\big{(}\sqrt{\frac{\log n}{n}}\big{)}.$ By triangle inequality,